Displaying report 1-1 of 1.
Reports until 12:25, Friday 09 August 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 12:25, Friday 09 August 2024 (11815)
OFC1 Power Budget

Attempting a power budget for loss measurement on OFC1. Haven't touched this cavity in a bit so some minor improvements/debugging were necessary:

-Moku 4 has issues with its wifi. I didn't care to debug this if someone else would like to try and fix this. For now I have a work around of plugging an Ethernet in and downloading the moku app to the NUC. Can connect to Moku 4 like this but it is slower than interacting via Ipad.

-Swapped cables so that I scan the laser frequency on moku 4. No error signal is present. Debugging steps: cable for the 1550 EOM is not shorted. Tank circuit for 1550 EOM is not shorted. Its not the cable. Confirmed it is outputting a voltage. The cavity is indeed scanning on the laser, I see transmission peaks. The REFL dips are very small. Improved 1550 alignment by tweaking input mirrors and maximized on REFL PD. REFL dips improved but still not good. Keeping REFL PD DC coupled and locking the cavity the light canceling on REFL is about 3/8s its original value (80 mv -> 30 mv). This is poor compared to the other cavity.

-The error amplitude is 400-600 uVpp. This used to be an order of magnitude higher and I am not sure why yet.

Displaying report 1-1 of 1.